# On the uniqueness of compiling graphs under the parity transformation

### v2

@@ 4.4.2 Polynomial-time compilation on rectangular plaquette layouts
or in other words, a given optimization problem that is represented as a graph can be compiled on a rectangular plaquette layout if and only if the graph restricted to the set of edges which are contained in an element of the cycle space is a complete bipartite graph. Hence, this necessary and sufficient condition immediately allows for the application of well-established algorithms for solving this computational problem in polynomial time. For example, in order to compute the restriction of a given graph $H=(V,E)\in\mathcal{H}$ onto ${E_{\mathcal{L}}}$ one could first apply forest growing algorithms [33], which usually have a time complexity of $\mathcal{O}(\left|V\right|+\left|E\right|)$ and compute a spanning forest of $H$. If the number of components of the graph $H$ is one, then iteratively collecting the unique fundamental cycles generates the restricted graph $H|_{E_{\mathcal{L}}}$. Note that each fundamental cycle can be determined by adding a non tree-edge to the spanning tree and calculating the two unique paths which start from the endpoints of the non tree-edge, traversing through their parents and end in the first common node. The running time of this part of the algorithm is proportional to $\left|V\right|\left|E\right|-\left|V\right|^{2}+\left|V\right|$, since the number of non tree-edges equals the dimension of the cycle space which is $\left|E\right|-\left|V\right|+1$. As a second step, by using classical search algorithms such as breadth-first or depth-first [33], which are also being used in forest growing algorithms, two independent partitions (in case the previously calculated restricted graph $H|_{E_{\mathcal{L}}}$ is bipartite) in $\mathcal{O}(\left|V\right|+\left|E\right|)$ running time can be computed. Finally, checking whether each vertex in the first partition has as many neighbors as the second partition elements has, completes the whole algorithm and yields a total time complexity proportional to $\left|V\right|\left|E\right|-\left|V\right|^{2}+\left|V\right|$. The corresponding cycle basis of the complete bipartite graph which compiles to a rectangular plaquette layout is given at the beginning of the proof of Theorem 4.10. Note that this algorithm can easily be extended to non-connected graphs.
@@ Acknowledgements
The authors would like to thank Anette Messinger and Christoph Fleckenstein for the detailed review of the manuscript as well as their crucial feedback. This work was supported by the Austrian Research Promotion Agency (FFG Project No. FO999896208). This research was funded in whole, or in part, by the Austrian Science Fund (FWF) SFB BeyondC Project No. F7108-N38, through a START grant under Project No. Y1067-N27 and I 6011. For the purpose of open access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission. This project was funded within the QuantERA II Programme that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No. 101017733.

### v3

@@ 4.4.2 Polynomial-time compilation on rectangular plaquette layouts
or in other words, a given optimization problem that is represented as a graph can be compiled on a rectangular plaquette layout if and only if the graph restricted to the set of edges which are contained in an element of the cycle space is a complete bipartite graph. Hence, this necessary and sufficient condition immediately allows for the application of well-established algorithms for solving this computational problem in polynomial time. For example, in order to compute the restriction of a given graph $H=(V,E)\in\mathcal{H}$ onto $E_{\mathcal{L}}$} one could first apply forest growing algorithms [33], which usually have a time complexity of $\mathcal{O}(\left|V\right|+\left|E\right|)$ and compute a spanning forest of $H$. If the number of components of the graph $H$ is one, then iteratively collecting the unique fundamental cycles generates the restricted graph $H|_{E_{\mathcal{L}}}$. Note that each fundamental cycle can be determined by adding a non tree-edge to the spanning tree and calculating the two unique paths which start from the endpoints of the non tree-edge, traversing through their parents and end in the first common node. The running time of this part of the algorithm is proportional to $\left|V\right|\left|E\right|-\left|V\right|^{2}+\left|V\right|$, since the number of non tree-edges equals the dimension of the cycle space which is $\left|E\right|-\left|V\right|+1$. As a second step, by using classical search algorithms such as breadth-first or depth-first [33], which are also being used in forest growing algorithms, two independent partitions (in case the previously calculated restricted graph $H|_{E_{\mathcal{L}}}$ is bipartite) in $\mathcal{O}(\left|V\right|+\left|E\right|)$ running time can be computed. Finally, checking whether the number of neighbors of each vertex in the first partition equals the number of elements of the second partition, completes the whole algorithm and yields a total time complexity proportional to $\left|V\right|\left|E\right|-\left|V\right|^{2}+\left|V\right|$. The corresponding cycle basis of the complete bipartite graph which compiles to a rectangular plaquette layout is given at the beginning of the proof of Theorem 4.10. Note that this algorithm can easily be extended to non-connected graphs.
@@ Acknowledgements
The authors would like to thank Anette Messinger and Christoph Fleckenstein for the detailed review of the manuscript as well as their crucial feedback. This work was supported by the Austrian Research Promotion Agency (FFG Project No. FO999896208). This research was funded in whole, or in part, by the Austrian Science Fund (FWF) SFB BeyondC Project No. F7108-N38 (DOI: 10.55776/F71), through a START grant under Project No. Y1067-N27 (DOI: 10.55776/Y1067) and I 6011 (DOI: 10.55776/I6011). For the purpose of open access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission. This project was funded within the QuantERA II Programme that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No. 101017733 (DOI: 10.3030/101017733).